I have a problem I can't complete, PLEASE HELP..... test problem
Find equation of a line
hERE IT IS: M=2 AND THROUGH THE POINT (-2,5) HERE'S WHAT i'VE DONE...
M=2 (-2,5) 2x2=Y2-5x2 4=Y2-5 so Y=2x+9
__________ +5 +5
2=y2-5 ________
______ 2
0-(-2) 9=Y2

Answer by stanbon(75887)

M=2 AND THROUGH THE POINT (-2,5)
--------------------------------------
There are two methods:
1st: Point-Slope Method
y-y1 = m(x-x1)
y-9
5 = 2(x--2)
y-5 = 2x+4
y = 2x + 9
-----------------
2nd: Slope-Intercept Method
y = mx + b
5 = 2*-2 + b
5 = -4 + b
b = 9
y = 2x+9
============
Cheers,
Stan H.